INAF queue observing, 2015 Apr 28-29

Observer: F. Cusano, R. Carini
Telescope Operator: David Gonzales Huerta
Support Astronomer: Dave Thompson on call
Instrument: LUCI

Summary

At the sunset the sky is mostly clear.
  • 02:50 sky mostly cloudy
  • 03:15 sky mostly clear again

Night duration (12\xB0 el twilight) 20:00-04:30 (3:00-11.30 UT) = 8 hrs 30 m

Observing time: 7.30 hrs

Weather time loss: 1.0 hrs (0%)

Data Summary

At the sunset the sky is mostly clear.
  • 02:50 sky mostly cloudy
  • 03:15 sky mostly clear again, humidity 67% , seeing 1.3

Project_RUN Objects Filter ExpTime Notes
Della Ceca D standard HIP56736_A0V HKspec    
Della Ceca D 2XMMJ100038.9+050955 HKspec 50 m 40 m good
Della Ceca E 2XMMJ123204.9+215254 K 2 h COMPLETED
Della Ceca E standard HIP62194_A0V K    
Rossi E GRB070521 HKspec 1 h 28 m seeing 1.5-2
Rossi E standard HR 6013 HKspec    

Details

03:15 begin with Della Ceca program, target D.
  • acquire the telluric star executeLUCIScript.sh Target_11D_tell.acq luci.20150429.0002-0004
  • executeLUCIScript.sh Target_11D_tell.spec luci.20150429.0005-0006

03:23 acquisition of science
  • executeLUCIScripts.sh Target_11D.acq luci.20150429.0007-0011
  • executeLUCIScripts.sh Target_11D.spec luci.20150429.0012-0016
  • seeing 1.25
  • 04:12 seeing 0.9-1.1
  • 04:40 many cirrus in the sky, image 0016 is not good
  • 04:45 we abort the acquisition, the spectrum is too faint

4:50 TO closes the dome for possible rain.

05:45 TO opens the dome again, we slew to the target E of Della Ceca program. the target D is out of spec.
  • executeLUCIScript.sh Target_11E.acq luci.20150429.0017-0020
  • executeLUCIScript.sh Target_11E.spec luci.20150429.0021-0032
  • at the beginning of the observation seeing 1.1, during the observation seeing variable between 1 and 1.5

08:23 slew to standard star
  • executeLUCIScript.sh Target_11E_tell.acq
  • problem with the guiding, TO refocused
  • 08:34 executeLUCIScript.sh Target_11E_tell.acq already problems , TO tries to solve them
  • 08:41 executeLUCIScript.sh Target_11E_tell.acq luci.20150429.0033-0034
  • 08:45 executeLUCIScript.sh Target_11E_tell.spec luci.20150429.0035 have stopped, the source now is not in the slit, we have to do the acquisitions one more time without guide stars, seeing is 2.7

08:50 we run the acquisition manually because we had some problems with the guide stars
  • acquisition Target_11E_tell.acq luci.20150429.0036-0046
  • acquisition Target_11E_tell.spec luci.20150429.0047-0048

09.06 slew to the Rossi Program Target E *executeLUCIScript.sh GRB070521_ls.acq luci.20150429.0049-0054
  • seeing 1.8
  • executeLUCIScript.sh GRB070521_ls_HK.spec luci.20150429.0055-0070
  • seeing 2.0
  • 10:00 seeing very variable 1.5-1.8

11:14 we slew to standard
  • executeLUCIScript.sh GRB070521telluric_ls.acq luci.20150429.0071-0073
  • seeing 2-2.5
  • executeLUCIScript.sh GRB070521telluric_ls_HK.spec luci.20150429.0074-0075

-- %USERSIG{RobertaCarini - 2015-04-29}%

Comments

 
Topic revision: r4 - 29 Apr 2015, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback